# How Malicious AI Swarms Can Threaten Democracy

### v3

@@
A step toward global coordination could be a distributed ‘AI Influence Observatory’ ecosystem — a network of academic groups, NGOs, and multilateral institutions. Its goal would be to standardize evidence, improve situational awareness, and enable faster collective response rather than impose top-down reputational penalties. To be practical, the ecosystem should rely on narrowly defined, privacy-preserving inputs and provide vetted researcher sandboxes for independent analysis. Civil-society reporting, investigative journalism, and whistleblower channels would complement technical signals, enabling triangulation across diverse evidence streams. For severe cross-border incidents, an impartial multilateral investigatory mechanism could evaluate claims and publish verified incident reports. The observatory’s verified incident reports could then serve as an impartial evidence base, enabling national or regional regulators to more effectively apply their own enforcement actions and economic sanctions.
@@
Because regulation and voluntary compliance face significant political resistance, a pragmatic approach targets underlying economic drivers. A key priority here is disrupting the commercial market that underpins large-scale manipulation, where private sellers offer services ranging from boosting vanity metrics to executing coordinated influence operations at remarkably low costs. Beyond detection, commercial-incentive levers can reduce profits from manipulation by domestic and foreign operators. Policies that may be helpful include adopting no-revenue policies for malicious swarm-proliferated content, discounting synthetic engagement in ranking and revenue-sharing, and publishing audited bot-traffic metrics. Safeguards must cover parties, campaigns, and officeholders, including party-linked media and contractors.
@@
Finally, companies should be required to promptly disclose when an account is flagged for behavior indicative of coordinated inauthentic activity, ensuring transparency while allowing for processes to address potential false positives. Policymakers should encourage—and where appropriate, incentivize—platforms to provide meaningful, privacy-preserving access for independent researchers so research can keep pace with evolving threats. At the same time, pre-bunking campaigns can help build cognitive resilience by empowering people and systems (“model immunization”) to spot the fingerprints of AI swarms. To strengthen structural defenses, interoperable “pro-user media” architectures, defined by empowering design principles that prioritize user well-being and epistemic health over maximizing viral engagement, can promote healthier information flows. At the same time, governments and technology firms should prioritize AI-safety research and fund independent measurement of misuse and societal impact. Taken together, these measures offer a layered strategy: immediate transparency to restore trust, proactive education to bolster citizens, resilient infrastructures to reduce systemic vulnerabilities, and sustained investment to monitor and adapt over time.

### v4

@@
To counter the speed, scale, independence, and adaptability of AI swarms, a step toward global coordination could be a distributed ‘AI Influence Observatory’ ecosystem – a network of academic groups, NGOs, and multilateral institutions. Its goal would be to standardize evidence, improve situational awareness, and enable faster collective response rather than impose top-down reputational penalties. To be practical, the ecosystem should rely on narrowly defined, privacy-preserving inputs and provide vetted researcher sandboxes for independent analysis. Civil-society reporting, investigative journalism, and whistleblower channels would complement technical signals, enabling triangulation across diverse evidence streams. For severe cross-border incidents, an impartial multilateral investigatory mechanism could evaluate claims and publish verified incident reports. The observatory’s verified incident reports could then serve as an impartial evidence base, enabling national or regional regulators to more effectively apply their own enforcement actions and economic sanctions.
@@
Because regulation and voluntary compliance face considerable political resistance, and because AI swarms make sophisticated manipulation cheaper and more effective, a pragmatic approach should target underlying economic drivers. A key priority here would be to disrupt the commercial market that underpins large-scale manipulation, where private sellers offer services ranging from boosting vanity metrics to executing coordinated influence operations at remarkably low costs. Beyond detection, commercial-incentive levers can reduce profits from manipulation by domestic and foreign operators. Policies that may be helpful include adopting no-revenue policies for malicious swarm-proliferated content, discounting synthetic engagement in ranking and revenue-sharing, and publishing audited bot-traffic metrics. Safeguards must cover parties, campaigns, and officeholders, including party-linked media and contractors.
@@
Finally, companies should be required to promptly disclose when an account is flagged for behavior indicative of coordinated inauthentic activity, ensuring transparency while allowing for processes to address potential false positives. Policymakers should encourage – and where appropriate, incentivize – platforms to provide meaningful, privacy-preserving access for independent researchers so research can keep pace with evolving threats. At the same time, pre-bunking campaigns can help build cognitive resilience by empowering people and systems (“model immunization”) to spot the fingerprints of AI swarms. To strengthen structural defenses, interoperable “pro-user media” architectures, defined by empowering design principles that prioritize user well-being and epistemic health over maximizing viral engagement, can promote healthier information flows. At the same time, governments and technology firms should prioritize AI-safety research and fund independent measurement of misuse and societal impact. Taken together, these measures offer a layered strategy: immediate transparency to restore trust, proactive education to bolster citizens, resilient infrastructures to reduce systemic vulnerabilities, and sustained investment to monitor and adapt over time.